what is the area of the regular 20​-gon with radius 9 mm?
Answer by Alan3354(69443)

what is the area of the regular 20-gon with radius 9 mm?
-------------
Divide it into 40 right triangles.
Each triangle has a base of 9 mm.
Each central angle is 9 degs.
tan(9) = h/b = h/9 --> h = b*tan(9)
-------
The area of each triangle is b*h/2
Find the area.
Multiply it by 40